# Agentic Large Language Models, a survey

### v2

@@ 4.1 Social Capabilities of LLMs
4.1.1 Conversation
As discussed in Section 1.3, the key advancement of instruction-tuned LLMs is their ability to interact using natural language. This requires a degree of formal linguistic competence, producing correct, grammatical sentences. However, the key factors for smooth and satisfying interactions are functional and pragmatic competence: the ability to understand what a user means and wants in a specific context (Mahowald et al., 2023). Various forms of finetuning improve the functional and pragmatic competence of LLMs (Ruis et al., 2023). Model size is also an important factor. However, the variation between different domains of functional and pragmatic understanding in LLMs is still large, and scores are overall below human performance (Sravanthi et al., 2024). One factor is that LLMs have less access to contextual information: they cannot see, hear, and otherwise sense the same as their human counterpart, nor do they have knowledge of previous interactions (Bender et al., 2021). With the shift to agentic LLMs this situation is improving, as they become equipped with memories, multi-modal capacities, and other tools that ground them in interactive contexts.
@@ 4.1.1 Conversation
Social etiquette and politeness in human-machine interaction have been studied for decades, see the review by Ribino (2023). Studies found that humans trust polite machines better when they adhere to social etiquette (Miller, 2005). Polite interactions lead to acceptance of machines as social entities, improving task performance and satisfaction (Miyamoto et al., 2021). LLM-chatbots are experienced as polite by users and, reversely, politeness of the user can drive the quality of the LLM output (Yin et al., 2024).
LLMs can detect affective and emotional states in language utterances (Broekens et al., 2023) and factor such information in their interaction behavior, becoming a more empathetic conversation partner (Yang et al., 2024d; Yan et al., 2024). For traditional LLMs, such empathy is limited to immediate conversational contexts. LLMs with access to additional contextual information or memory have further improved empathetic abilities (Sravanthi et al., 2024).
4.1.2 Strategic Behavior
Game theory is the field that studies strategic behavior by agents (Von Neumann and Morgenstern, 2007). The field studies strategic questions of allocation of scarce resources, fairness, and social dilemmas (Jones, 2000). There is a long history in this field of using machine learning (Fatima et al., 2024). Recently, researchers have studied how LLM behavior differs from that of other types of computational architectures as well as from humans. In this section we discuss work on unenhanced, non-agentic models that are given a prompt or script to take part in a social experiment or game.
@@ 4.1.2 Strategic Behavior
Perhaps the best-known social dilemma is the Prisoner’s Dilemma (Rapoport, 1965; Axelrod, 1980; Poundstone, 2011). One study models the Iterated Prisoner’s Dilemma (Fontana et al., 2024) in LLaMa2, LLaMa3, and GPT3.5. They find that models are cautious, favoring cooperation over defection only when the opponent’s defection rate is low. Overall, LLMs behave at least as cooperatively as the typical human player, although there are substantial differences among models. In particular, LLaMa2 and GPT3.5 are more cooperative than humans, and especially forgiving and non-retaliatory for opponent defection rates below 30%. More similar to humans, LLaMa3 exhibits consistently uncooperative and exploitative behavior unless the opponent always cooperates.
@@
Akata et al. (2023) set up different LLMs to play various repeated games (GPT-3, GPT-3.5, and GPT-4). The LLMs are particularly good at games where valuing their own self-interest pays off, such as the iterated Prisoner’s Dilemma. However, they are less good in games that require coordination, such as Battle of the Sexes. GPT-4’s behavior is shown to be sensitive to additional information provided about the other player, as well as prompts asking it to predict the other player’s actions before making a choice. This effect is studied further by Lorè and Heydari (2023), who distinguish between abstract strategic reasoning (needed to determine an optimal strategy given the structure of a game) and responsiveness to contextual framing (such as you are dealing with a diplomatic relation or a casual friend). They find that abstract reasoning capacity is highest in LLaMa-2, followed by GPT-4. GPT-3.5 shows little abstract reasoning capacity and is highly sensitive to contextual framing. The picture that emerges from these initial studies is that LLMs have varied strategic proficiencies in economic games, and that they can relatively easily be influenced by additional information in the prompt.
@@
Recent systematic benchmarking has corroborated these results. GTBench (Game Theory benchmark) (Duan et al., 2024) covers Tic-Tac-Toe, Connect-4, Kuhn Poker, Breakthrough, Liar’s Dice, Blind Auction, Negotiation, Nim, Pig, and the Iterated Prisoner’s Dilemma. They find that LLMs fail in complete and deterministic games yet are competitive in probabilistic gaming scenarios; most open-source LLMs (such as LLaMa) are less competitive than commercial LLMs (GPT-4) in complex games (except for LLaMa-3-70b-Instruct, which does perform well). In addition, code-pretraining greatly benefits strategic reasoning, while advanced reasoning methods such as Chain of Thought and Tree of Thoughts do not always help.
@@
EgoSocialArena (Hou et al., 2024) focuses on cognitive, situational, and behavioral intelligence, see Figure 19. All tested models (including OpenAI o1-preview) lag 11% behind humans. The superiority of o1-preview is mainly attributed to its logical reasoning and mathematical abilities that find deep patterns in the data. Comparing the performance of a small version of LLaMa (LLaMA3-8B-Chat) with a large version (LLaMA3-70B-Chat), they find that model size does not significantly help improve social intelligence. In this study, LLMs show improved theory of mind reasoning ability when operating from a first-person perspective than from the third-person, providing counterweight to contrasting findings by (Kim et al., 2023).
4.1.3 Theory of Mind
@@
An advanced capability that enables social interaction in humans is theory of mind. Humans use theory of mind to attribute mental states to others and reason about the world from their perspective (Premack and Woodruff, 1978; Apperly, 2011). Theory of mind enables us to make social judgments and to plan future steps in interactions, since we can imagine someone else’s reaction. Theory of mind is related to planning (Section 2.1.3) and self reflection (Section 2.2.2) in LLMs.
@@
Early experiments by Kosinski (2023, 2024) showed that models could pass tests for assessing theory of mind in children and clinical populations. This led to the claim that theory of mind had spontaneously emerged in LLMs, given that they were neither designed nor trained specifically to perform theory of mind tasks. The experiments were criticized due to the occurrence of false-belief test questions (and correct answers) in the training data (Ullman, 2023; Shapira et al., 2024). Recently a more nuanced perspective formed, as specific theory of mind benchmarks were introduced (Kim et al., 2023; Chen et al., 2024b; Wang et al., 2024a), other modalities were integrated (van Berkel, 2024; Strachan et al., 2024), integrations with older model architectures were explored (Jin et al., 2024a), and direct comparisons to human performance were made (Van Dijk et al., 2023; Strachan et al., 2024).
@@
An application domain of theory of mind is social judgment. LLMs have been shown to outperform average human scores on a social-situational judgment task (Mittelstädt et al., 2024). Results from five different LLM-based chatbots were compared with responses of 276 human participants, showing that Claude, Copilot and You.com’s smart assistant performed significantly better than human subjects at proposing suitable behaviors in the descriptions of social situations. Moreover, their options for different behavior aligned well with expert ratings.
@@
Although the results of early experiments on the emergence of theory of mind in LLMs were less convincing, stronger commercial LLMs are steadily improving, scoring at or sometimes above average human level on standardized tests. Further research and discussion are needed to show whether high scores on such tests mean that LLMs have generalizable forms of theory of mind (Goldstein and Levinstein, 2024; Hu et al., 2025; van der Meulen et al., 2025).
4.2 Role-Based Interaction
@@
LLMs are being used in the fields of multi-agent systems and agent-based simulation (Gao et al., 2024), which have a long research tradition (Epstein and Axtell, 1996; Macal and North, 2010). Multi-agent approaches simulate individual agents and their interactions in an environment that is often virtual, but can also be physical (Steels, 1995; Shoham and Leyton-Brown, 2008). Complex dynamics can emerge between agents with basic perceptive, reasoning, and decision-making abilities. Agent-based approaches are often used as a bridge between theoretical and empirical work, allowing for exploration and hypothesis testing in domains where working with human agents is unethical, costly, or otherwise difficult.
@@
Challenges in modeling realistic agent behavior, as well as the computational cost of simulating multi-agent societies, have often impeded realistic multi-agent experiments. Advances in agentic LLMs and computational infrastructure for multi agent simulations (Rutherford et al., 2024) are changing this situation, and have given an impulse to research in experimental computational game theory. Creating agents that use LLMs has enabled researchers to overcome existing limitations, by letting agents communicate in natural language. This allowed for the exploration of new territory in the domains of game theory, role-based interactions, and team work.
4.2.1 Strategic Behavior in Multi-LLM Environments
@@
The MAgIC study (Xu et al., 2024a) uses social deduction games (Undercover and Chameleon) and game theoretic scenarios such as Cost Sharing, Multi player Prisoner’s Dilemma, and Public Good. From these games, seven features are extracted: Rationality, Judgement, Reasoning, Deception, Self-awareness, Cooperation, Coordination, as shown in Figure 20. LLMs are evaluated on these critical abilities in multi-agent environments. GPT-o1 and GPT-4 score significantly better than the other LLMs. Interestingly, LLMs score generally high on Judgement, Rationality and Cooperation, but some also on Deception. Further, they all score lower on Reasoning and all but one score low on Coordination. The exception here is GPT-o1 enhanced with probabilistic graphic modeling, an implementation of a theory of mind-like competence inspired by Koller and Friedman (2009). The authors show that probabilistic graphic modeling boosts LLM performance on their evaluation-games across the board. This fits with the generally accepted idea that humans rely on their theory of mind in game-theoretical scenarios.
GAMA-Bench is a benchmark for multi agent games (Huang et al., 2024a) that covers Guess 2/3 of the Average, El Farol Bar, Divide the Dollar, Public Goods Game, Diner’s Dilemma, Sealed-bid Auction, Battle Royale, and Pirate Game. The results show that while GPT-3.5 is robust, its generalizability is limited. Here, performance can be improved through approaches such as Chain of Thought. Additionally, evaluations across various LLMs were conducted, showing that GPT-4 outperforms other models. Moreover, increasingly higher scores across three iterations of GPT-3.5 demonstrate marked advancements in the model’s intelligence with each update.
Alympics is a platform for complex strategic multi agent gaming problems (Mao et al., 2023). It provides a controlled playground for simulating human-like strategic interactions with LLM-driven agents. Figure 21 shows an example of their water allocation challenge, a complex strategy game in which scarce resources for survival must be distributed across multiple rounds.
AucArena simulates auctions, on LLaMa 2.13b, Mistral 7b, Mixtral 8x7b, Gemini 1.0, and GPT 3.5 and 4.0 (Chen et al., 2023a). The authors find that LLMs such as GPT-4 possess important skills for auction participation, such as budget management and goal-focus. However, they also find that performance varies, pointing to opportunities for improvement.
4.2.2 Role-Based Task Solving and Team Work
@@
In the CAMEL framework (Communicative Agents for “Mind” Exploration) (Li et al., 2023a), two LLMs have the predefined roles to perform, for example, a coding task (See Figure 22). They cooperatively drive a conversation without continuous human prompting. By using inception prompting and role descriptions, the agents stay in character and collaborate toward the goal by breaking down complex problems in manageable steps through dialogue. During each interaction step the LLM agents effectively generate their own inference-time training data, making the cognitive process visible for human inspection while enhancing autonomous task performance.
@@
Other studies have paired an LLM creator or generator with an LLM judge or critic. In this way the generative power of LLMs is leveraged, while adherence to rules or quality guidelines is enforced. Constitutional AI (Bai et al., 2022) employs one LLM to critique another LLM’s responses against a set of ethical or quality guidelines, and to suggest revisions. The authors show that this kind of two-agent feedback loop yields refined final outputs that is aligned with desired principles.
@@
Another form of role-based interaction is the use of debate or discussion between LLMs to improve reasoning and task performance. Du et al. (2024) demonstrated that when multiple LLM instances propose answers and critique each other’s reasoning through several rounds of debate, they can reach a more accurate consensus answer with higher factual correctness. This approach, described as a society of minds, significantly reduced reasoning errors and hallucinations in tasks like math word problems and factual QA.
@@
Similarly, Chan et al. (2024) propose a Multi-Agent Debate (MAD) setup where two LLM agents take opposing sides in a tit-for-tat debate while a third agent acts as a judge. The role of the judge is to guide the discussion towards a final solution. The structured debates encouraged divergent thinking and could even push a weaker model (such as GPT-3.5) to outperform a stronger model (such as GPT-4) on certain challenging problems by combining strengths of each agent.
@@
Motivated by Minsky’s society of minds (Minsky, 1988), a multi-agent framework has been designed as a round table conference among diverse LLM agents (Chen et al., 2023b). The framework enhances collaborative reasoning between LLM agents via multiple rounds of discussion. The agents should learn to convince other agents to improve their answers. Experiments on seven benchmarks demonstrate that a confidence-weighted voting mechanism significantly improves LLMs’ reasoning. Furthermore, the authors find that diversity (different models) is critical for performance. Again inspired by Minsky, MindStorms introduces an LLM-based implementation (Zhuge et al., 2023) on the CAMEL framework (Li et al., 2023a). Extensive experiments are reported with up to 129 agents solving common AI problems: visual question answering, image captioning, text-to-image synthesis, 3D generation, egocentric retrieval, embodied AI, and general language-based task solving. They found that, in specific applications, mindstorms among many members outperform those among fewer members, and longer mindstorms outperform shorter ones.
@@
Related to debate and discussion setups, researchers have explored teacher-learner dynamics with LLMs, where an expert LLM provides hints or feedback to a less capable LLM on a task, mirroring human tutoring (Zhou et al., 2024c). These role-alignments leverage the idea that one agent’s knowledge or oversight can correct the other’s mistakes, leading to more robust performance. AutoGen (Wu et al., 2023) is designed to facilitate the development of multi agent LLM applications that span a broad spectrum of domains and complexities. The programming paradigm is centered around agent-agent conversations. Experiments demonstrate the effectiveness of the framework in example applications ranging from mathematics, coding, question answering, operations research, online decision-making, to entertainment.
@@
ChatEval is a multi-agent system to improve text summarization (Chan et al., 2023). Noting that the quality of human text summarization improves when multiple annotators collaborate, the authors created a multi-agent debate framework, moving beyond single-agent prompting strategies, including debater agents, diverse role specification, and different communication strategies (see Figure 23).
@@
Sotopia is another role-playing environment for multi-agent interaction (Zhou et al., 2023b). In Sotopia, agents coordinate, collaborate, exchange, and compete with each other to achieve complex social goals. In experiments with LLM-agents and humans, GPT-4 achieves a significantly lower goal completion rate than humans and struggles to exhibit social commonsense reasoning and strategic communication skills. The contrast between GPT-4’s lower performance in Sotopia and good performance on other metrics of social reasoning (see Section 4.1.3) is most likely explained by Sotopia’s focus on strategizing and goal-directedness, aspects on which GPT-4 is known to score lower (Hou et al., 2024).
@@
To simulate strategic interaction and cooperative decision-making, researchers have introduced GovSim (Piatti et al., 2024). They study how ethical considerations, strategic planning, and negotiation skills impact cooperative outcomes. Most LLMs fail to achieve an equilibrium since they fail to understand the long-term effects of their actions on the group. GPT-4o performed best. Interestingly, the introduction of a special universalization reasoning language (Levine et al., 2020) (prompting models to ask the Kantian question: What if everybody does that?) allowed more models to achieve a sustainable outcome. Related results were demonstrated in NegotiationArena, introduced by Bianchi et al. (2024). They showed how LLM agents can conduct complex negotiations through flexible dialogue in negotiation settings. The flexible dialogues significantly improved negotiation outcomes by employing different behavioral strategies.
@@
Social interaction in an extreme setting was studied in (Campedelli et al., 2024). Inspired by the Stanford Prison experiment (Zimbardo, 1972), the emergence of persuasive and abusive behavior is studied in a setting of prisoners versus prison guards. It was found that the assigned personality of prisoner and guard impact both persuasiveness and the emergence of anti-social behavior. Anti-social behavior emerged by simply assigning the agent’s roles, which is a parallel to the original experiments involving human participants.
4.3 Simulating Open-ended Societies

### v3

@@ 4.1. Social Capabilities of LLMs
4.1.1. Conversation
As discussed in Section 1.3, the key advancement of instruction-tuned LLMs is their ability to interact using natural language. This requires a degree of formal linguistic competence, producing correct, grammatical sentences. However, the key factors for smooth and satisfying interactions are functional and pragmatic competence: the ability to understand what a user means and wants in a specific context (mahowald2023dissociating). Various forms of finetuning improve the functional and pragmatic competence of LLMs (ruis2023goldilocks). Model size is also an important factor. However, the variation between different domains of functional and pragmatic understanding in LLMs is still large, and scores are overall below human performance (sravanthi2024-pub). One factor is that traditional LLMs have less access to contextual information: they cannot see, hear, and otherwise sense the same as their human counterpart, nor do they have knowledge of previous interactions (bender2021stochastic). With the shift to agentic and multi-modal LLMs this situation is improving, as they become equipped with memories, multi-modal capacities, and other tools that ground them in interactive contexts.
@@ 4.1.1. Conversation
Social etiquette and politeness in human-machine interaction have been studied for decades, see the review by Ribino2023role. Studies found that humans trust polite machines better when they adhere to social etiquette (Miller2005). Polite interactions lead to acceptance of machines as social entities, improving task performance and satisfaction (Miyamoto2021). LLM-chatbots are experienced as polite by users and, reversely, politeness of the user can drive the quality of the LLM output (yin-etal-2024-respect).
LLMs can detect affective and emotional states in language utterances (broekens2023fine) and factor such information in their interaction behavior, becoming a more empathetic conversation partner (yang2024enhancing; yan2024talkhumanlikeagentsempathetic). For traditional LLMs, such empathy is limited to immediate conversational contexts. LLMs with access to additional contextual information or memory have further improved empathetic abilities (sravanthi2024-pub).
4.1.2. Strategic Behavior
Game theory is the field that studies strategic behavior by agents (von2007theory). The field studies strategic questions of allocation of scarce resources, fairness, and social dilemmas (jones2000game). There is a long history in this field of using machine learning (fatima2024learning). Recently, researchers have studied how LLM behavior differs from that of other types of computational architectures as well as from humans. In this section we discuss work on unenhanced, non-agentic models that are given a prompt or script to take part in a social experiment or game.
@@ 4.1.2. Strategic Behavior
Perhaps the best-known social dilemma is the Prisoner’s Dilemma (rapoport1965prisoner; axelrod1980effective; poundstone2011prisoner). A study by fontana2024nicer models the iterated Prisoner’s Dilemma in LLaMa2, LLaMa3, and GPT3.5. They find that models are cautious, favoring cooperation over defection only when the opponent’s defection rate is low. Overall, LLMs behave at least as cooperatively as the typical human player, although there are substantial differences among models. In particular, LLaMa2 and GPT3.5 are more cooperative than humans, and especially forgiving and non-retaliatory for opponent defection rates below 30%. More similar to humans, LLaMa3 exhibits consistently uncooperative and exploitative behavior unless the opponent always cooperates.
@@
akata2023playing set up different LLMs to play various repeated games (GPT-3, GPT-3.5, and GPT-4). The LLMs are particularly good at games where valuing their own self-interest pays off, such as the iterated Prisoner’s Dilemma. However, they are less good in games that require coordination, such as Battle of the Sexes. GPT-4’s behavior is shown to be sensitive to additional information provided about the other player, as well as prompts asking it to predict the other player’s actions before making a choice. This effect is studied further by lore2023strategic, who distinguish between abstract strategic reasoning (needed to determine an optimal strategy given the structure of a game) and responsiveness to contextual framing (such as you are dealing with a diplomatic relation or a casual friend). They find that abstract reasoning capacity is highest in LLaMa-2, followed by GPT-4. GPT-3.5 shows little abstract reasoning capacity and is highly sensitive to contextual framing. The picture that emerges from these initial studies is that LLMs have varied strategic proficiencies in economic games, and that they can relatively easily be influenced by additional information in the prompt.
@@
Recent systematic benchmarking has corroborated these results. GTBench (Game Theory benchmark) (duan2024gtbench) covers Tic-Tac-Toe, Connect-4, Kuhn Poker, Breakthrough, Liar’s Dice, Blind Auction, Negotiation, Nim, Pig, and the Iterated Prisoner’s Dilemma. They find that LLMs fail in complete and deterministic games yet are competitive in probabilistic gaming scenarios; most open-source LLMs (such as LLaMa) are less competitive than commercial LLMs (GPT-4) in complex games (except for LLaMa-3-70b-Instruct, which does perform well). In addition, code-pretraining greatly benefits strategic reasoning, while advanced reasoning methods such as Chain of Thought and Tree of Thoughts do not always help.
@@
EgoSocialArena (hou2024entering) focuses on cognitive, situational, and behavioral intelligence, see Figure 19. All tested models (including OpenAI o1-preview) lag 11% behind humans. The superiority of o1-preview is mainly attributed to its logical reasoning and mathematical abilities that find deep patterns in the data. Comparing the performance of a small version of LLaMa (LLaMA3-8B-Chat) with a large version (LLaMA3-70B-Chat), they find that model size does not significantly help improve social intelligence. In this study, LLMs show improved theory of mind reasoning ability when operating from a first-person perspective than from the third-person, providing counterweight to contrasting findings by (kim-etal-2023-fantom).
4.1.3. Theory of Mind
@@
An advanced capability that enables social interaction in humans is theory of mind. Humans use theory of mind to attribute mental states to others and reason about the world from their perspective (premack1978does; apperly2011mindreaders). Theory of mind enables us to make social judgments and to plan future steps in interactions, since we can imagine someone else’s reaction. Theory of mind is related to planning (Section 2.1.3) and self reflection (Section 2.2.2) in LLMs.
@@
Early experiments by kosinski2023theory; kosinski2024evaluating showed that models could pass tests for assessing theory of mind in children and clinical populations. This led to the claim that theory of mind had spontaneously emerged in LLMs, given that they were neither designed nor trained specifically to perform theory of mind tasks. The experiments were criticized due to the occurrence of false-belief test questions (and correct answers) in the training data (ullman2023large; shapira2023clever). Recently a more nuanced perspective formed, as specific theory of mind benchmarks were introduced (kim-etal-2023-fantom; chen2024tombenchbenchmarkingtheorymind; wang2024tmgbenchsystematicgamebenchmark), other modalities were integrated (razothesis2024; strachan2024gpt4oreadsmindeyes), integrations with older model architectures were explored (jin2024mmtomqamultimodaltheorymind), and direct comparisons to human performance were made (van2023large; strachan2024gpt4oreadsmindeyes).
@@
An application domain of theory of mind is social judgment. LLMs have been shown to outperform average human scores on a social-situational judgment task (mittelstadt2024large). Results from five different LLM-based chatbots were compared with responses of 276 human participants, showing that Claude, Copilot and You.com’s smart assistant performed significantly better than human subjects at proposing suitable behaviors in the descriptions of social situations. Moreover, their options for different behavior aligned well with expert ratings.
@@
Although the results of early experiments on the emergence of theory of mind in LLMs were less convincing, stronger commercial LLMs are steadily improving, scoring at or sometimes above average human level on standardized tests. Further research and discussion are needed to show whether high scores on such tests mean that LLMs have generalizable forms of theory of mind (goldstein2024doesChatGPTmind; hu2025reevaluatingtheorymindevaluation; vandermeulen2025properlyimplementingtheorymind).
4.2. Role-Based Interaction
@@
LLMs are being used in the fields of multi-agent systems and agent-based simulation (gao2024large), which have a long research tradition (epstein1996growing; macal_tutorial_2010). Multi-agent approaches simulate individual agents and their interactions in an environment that is often virtual, but can also be physical (steels1995self; shoham2008multiagent). Complex dynamics can emerge between agents with basic perceptive, reasoning, and decision-making abilities. Agent-based approaches are often used as a bridge between theoretical and empirical work, allowing for exploration and hypothesis testing in domains where working with human agents is unethical, costly, or otherwise difficult.
@@
Challenges in modeling realistic agent behavior, as well as the computational cost of simulating multi-agent societies, have often impeded realistic multi-agent experiments. Advances in agentic LLMs and computational infrastructure for multi agent simulations (rutherford2024jaxmarl) are changing this situation, and have given an impulse to research in experimental computational game theory. Creating agents that use LLMs has enabled researchers to overcome existing limitations, by letting agents communicate in natural language. This allowed for the exploration of new territory in the domains of game theory, role-based interactions, and team work.
4.2.1. Strategic Behavior in Multi-LLM Environments
@@
The MAgIC study (xu2024magic) uses social deduction games (Undercover and Chameleon) and game theoretic scenarios such as Cost Sharing, Multi player Prisoner’s Dilemma, and Public Good. From these games, seven features are extracted: Rationality, Judgement, Reasoning, Deception, Self-awareness, Cooperation, Coordination, as shown in Figure 20. LLMs are evaluated on these critical abilities in multi-agent environments. GPT-o1 and GPT-4 score significantly better than the other LLMs. Interestingly, LLMs score generally high on Judgement, Rationality and Cooperation, but some also on Deception. Further, they all score lower on Reasoning and all but one score low on Coordination. The exception here is GPT-o1 enhanced with probabilistic graphic modeling, an implementation of a theory of mind-like competence inspired by kollerfriendman2009pgm. The authors show that probabilistic graphic modeling boosts LLM performance on their evaluation-games across the board. This fits with the generally accepted idea that humans rely on their theory of mind in game-theoretical scenarios.
GAMA-Bench is a benchmark for multi agent games (huang2024far) that covers Guess 2/3 of the Average, El Farol Bar, Divide the Dollar, Public Goods Game, Diner’s Dilemma, Sealed-bid Auction, Battle Royale, and Pirate Game. The results show that while GPT-3.5 is robust, its generalizability is limited. Here, performance can be improved through approaches such as Chain of Thought. Additionally, evaluations across various LLMs were conducted, showing that GPT-4 outperforms other models. Moreover, increasingly higher scores across three iterations of GPT-3.5 demonstrate marked advancements in the model’s intelligence with each update.
Alympics is a platform for complex strategic multi agent gaming problems (mao2023alympics). It provides a controlled playground for simulating human-like strategic interactions with LLM-driven agents. Figure 21 shows an example of their water allocation challenge, a complex strategy game in which scarce resources for survival must be distributed across multiple rounds.
AucArena simulates auctions, on LLaMa 2.13b, Mistral 7b, Mixtral 8x7b, Gemini 1.0, and GPT 3.5 and 4.0 (chen2023put). The authors find that LLMs such as GPT-4 possess important skills for auction participation, such as budget management and goal-focus. However, they also find that performance varies, pointing to opportunities for improvement.
4.2.2. Role-Based Task Solving and Team Work
@@
In the CAMEL framework (Communicative Agents for “Mind” Exploration) (li2023camel), two LLMs have the predefined roles to perform, for example, a coding task (See Figure 22). They cooperatively drive a conversation without continuous human prompting. By using inception prompting and role descriptions, the agents stay in character and collaborate toward the goal by breaking down complex problems in manageable steps through dialogue. During each interaction step the LLM agents effectively generate their own inference-time training data, making the cognitive process visible for human inspection while enhancing autonomous task performance. We will discuss CAMEL furhter in Section 4.4.2
@@
Other studies have paired an LLM creator or generator with an LLM judge or critic. In this way the generative power of LLMs is leveraged, while adherence to rules or quality guidelines is enforced. Constitutional AI (bai2022constitutional) employs one LLM to critique another LLM’s responses against a set of ethical or quality guidelines, and to suggest revisions. The authors show that this kind of two-agent feedback loop yields refined final outputs that is aligned with desired principles.
@@
Another form of role-based interaction is the use of debate or discussion between LLMs to improve reasoning and task performance. du2024debate demonstrated that when multiple LLM instances propose answers and critique each other’s reasoning through several rounds of debate, they can reach a more accurate consensus answer with higher factual correctness. This approach, described as a society of minds, significantly reduced reasoning errors and hallucinations in tasks like math word problems and factual QA.
@@
Similarly, chan2024MAD propose a Multi-Agent Debate (MAD) setup where two LLM agents take opposing sides in a tit-for-tat debate while a third agent acts as a judge. The role of the judge is to guide the discussion towards a final solution. The structured debates encouraged divergent thinking and could even push a weaker model (such as GPT-3.5) to outperform a stronger model (such as GPT-4) on certain challenging problems by combining strengths of each agent.
@@
Motivated by Minsky’s society of minds (minsky1988society), a multi-agent framework has been designed as a round table conference among diverse LLM agents (chen2023reconcile). The framework enhances collaborative reasoning between LLM agents via multiple rounds of discussion. The agents should learn to convince other agents to improve their answers. Experiments on seven benchmarks demonstrate that a confidence-weighted voting mechanism significantly improves LLMs’ reasoning. Furthermore, the authors find that diversity (different models) is critical for performance. Again inspired by Minsky, MindStorms introduces an LLM-based implementation (zhuge2023mindstorms) on the CAMEL framework (li2023camel). Extensive experiments are reported with up to 129 agents solving common AI problems: visual question answering, image captioning, text-to-image synthesis, 3D generation, egocentric retrieval, embodied AI, and general language-based task solving. They found that, in specific applications, mindstorms among many members outperform those among fewer members, and longer mindstorms outperform shorter ones.
@@
Related to debate and discussion setups, researchers have explored teacher-learner dynamics with LLMs, where an expert LLM provides hints or feedback to a less capable LLM on a task, mirroring human tutoring (zhou2024sotopia). These role-alignments leverage the idea that one agent’s knowledge or oversight can correct the other’s mistakes, leading to more robust performance. AutoGen (wu2023autogen) is designed to facilitate the development of multi agent LLM applications that span a broad spectrum of domains and complexities. The programming paradigm is centered around agent-agent conversations. Experiments demonstrate the effectiveness of the framework in example applications ranging from mathematics, coding, question answering, operations research, online decision-making, to entertainment.
@@
ChatEval is a multi-agent system to improve text summarization (chan2023chateval). Noting that the quality of human text summarization improves when multiple annotators collaborate, the authors created a multi-agent debate framework, moving beyond single-agent prompting strategies, including debater agents, diverse role specification, and different communication strategies (see Figure 23).
@@
Sotopia is another role-playing environment for multi-agent interaction (zhou2023sotopia). In Sotopia, agents coordinate, collaborate, exchange, and compete with each other to achieve complex social goals. In experiments with LLM-agents and humans, GPT-4 achieves a significantly lower goal completion rate than humans and struggles to exhibit social commonsense reasoning and strategic communication skills. The contrast between GPT-4’s lower performance in Sotopia and good performance on other metrics of social reasoning (see Section 4.1.3) is most likely explained by Sotopia’s focus on strategizing and goal-directedness, aspects on which GPT-4 is known to score lower (hou2024entering).
@@
To simulate strategic interaction and cooperative decision-making, researchers have introduced GovSim (piatti2024cooperate). They study how ethical considerations, strategic planning, and negotiation skills impact cooperative outcomes. Most LLMs fail to achieve an equilibrium since they fail to understand the long-term effects of their actions on the group. GPT-4o performed best. Interestingly, the introduction of a special universalization reasoning language (levine2020logic) (prompting models to ask the Kantian question: What if everybody does that?) allowed more models to achieve a sustainable outcome. Related results were demonstrated in NegotiationArena, introduced by bianchi2024well. They showed how LLM agents can conduct complex negotiations through flexible dialogue in negotiation settings. The flexible dialogues significantly improved negotiation outcomes by employing different behavioral strategies.
@@
Social interaction in an extreme setting was studied in (campedelli2024want). Inspired by the Stanford Prison experiment (zimbardo1972stanford), the emergence of persuasive and abusive behavior is studied in a setting of prisoners versus prison guards. It was found that the assigned personality of prisoner and guard impact both persuasiveness and the emergence of anti-social behavior. Anti-social behavior emerged by simply assigning the agent’s roles, which is a parallel to the original experiments involving human participants.
4.3. Simulating Open-ended Societies